OpenStudy (anonymous):

PLEASE give me guidance? What is the value of y in the solution to the following system of equations?5x - 3y = -32x - 6y = -6 Answers: -1 0 1 2

OpenStudy (anonymous):

see we get y=37/21 according to your equations

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

You did something wrong with the question..check it

OpenStudy (anonymous):

5x - 3y = -32x - 6y = -6

OpenStudy (anonymous):

You break it in two?

OpenStudy (anonymous):

well, if you are right with your question, then y=37/21 but you do not have it in your answer choices?

OpenStudy (anonymous):

yes, break it into two as 5x-3y= -6 and -32x-6y = -6

OpenStudy (anonymous):

No, it has to be one of the answers I provided

OpenStudy (anonymous):

(The answer I got was -1)

OpenStudy (anonymous):

show how it is -1

OpenStudy (anonymous):

Thanks dude for your support @bhaskarbabu

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

I see how you would, but you have to break it into one digit

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Do it carefully, its 37/21

OpenStudy (anonymous):

Heh, rushed through that, yes y=37/21.

OpenStudy (anonymous):

1.7619047619

OpenStudy (anonymous):

Are you sure you aren't missing a piece of one of the equations?

OpenStudy (anonymous):

Does 5x-37 actually equal something else?

OpenStudy (anonymous):

yes,that's it... 1.762

OpenStudy (anonymous):

My apologies, but that's what it was written as

OpenStudy (anonymous):

Would it be a positive 1 instead...?

OpenStudy (anonymous):

It is more reasonable to be positive 2

OpenStudy (anonymous):

Not an option

OpenStudy (anonymous):

What about this one? Use the substitution method to solve the following system of equations.x + y = 32x + 2y = 10

OpenStudy (anonymous):

x= -1 and y = 21

OpenStudy (anonymous):

It's either 1 or -1

OpenStudy (anonymous):

So it's negative one?!!! I WAS RIGHT!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

/8, Assuming the second equation, -32x-6y=-6 is correct, and the solution for y must be /16, -in the set, {-1, 0, 1, 2}, then solving for possible values of x: \[x=\frac{3}{16}-\frac{3}{16}[-1, 0, 1, 2] = [\frac{3}{8}, \frac{3}{16}, 0, \frac{-3}{16}]\] And your first equation might actually be 5x-3y=[39/16, 15/16, -3, -111/16]

OpenStudy (anonymous):

Thank you for breaking it down for me, @CliffSedge

OpenStudy (anonymous):

I just noticed your picture is of you or someone on a cliff and your username has the word "cliff" in it haha

OpenStudy (anonymous):

So, all in all, -1 was correct?

OpenStudy (anonymous):

In other words, in the more-than-likely event that the first equation is actually 5x-3y=-3, the solution to the system would be (0,1)

OpenStudy (anonymous):

A positive 1?

OpenStudy (anonymous):

How did you get zero, sir?

OpenStudy (anonymous):

He/she corrected the equations you gave at the beginning of this question thread...then you get x=0 and y=1

OpenStudy (anonymous):

Thank you, my sincerest apologies....@AbhimanyuPudi

OpenStudy (anonymous):

The possible answer choices for y that you posted were {-1, 0, 1, 2}. I put all of those into the second equation and solved for x and got {3/8, 3/16, 0, -3/16} Putting all those solutions into the first equation, I got other possibilities for what 5x-3y could equal. All the other coefficients were integers, so I induced that out of that set, that -3 was most likely. In that case, the solution would be (0,1).

OpenStudy (anonymous):

Thank you both @AbhimanyuPudi and @CliffSedge

OpenStudy (anonymous):

OHHHHHHH!!!!

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

It's a weak conjecture, but seems more likely than the solution being 37/21.

OpenStudy (anonymous):

Agreed

OpenStudy (anonymous):

a strong assumption

OpenStudy (anonymous):

If the correct answer isn't in the list of options, then either the question or the list of possible answers is wrong. It's hard to tell which; there are so many possibilities.

OpenStudy (anonymous):

I have a strong theory here @InsanelyChaotic @AbhimanyuPudi 5x-3y=-32x-6y=-6, like the other question is a case of not putting a space or line between 5x-3y=-3 2x-6y=-6 If you solve that system, you'll see that (0,1) is correct.

OpenStudy (anonymous):

I'm sorry, you are all idiots, and especially @InsanelyChaotic yes, it was two separate equations, and was listed as two separate equations on two separate lines of text in the UNIT TEST!!!

OpenStudy (anonymous):

Thank you for figuring that out @CliffSedge *medal!*